OpenStudy (linn99123):

NEED HELP FAST WILL GIVE MEDAL AND FAN A circular piece of metal has a radius of 3.5 meters. The metal sells for $3.25 per square meter. What is the total cost of the circular piece of metal? A. $11.38 B. $38.47 C. $71.44 D. $125.01

OpenStudy (linn99123):

@jigglypuff314 can u help

jigglypuff314 (jigglypuff314):

area of circle = pi r^2 r = 3.5 so = pi (3.5 ^2) =

jigglypuff314 (jigglypuff314):

then multiply what you get by 3.25 :)

OpenStudy (linn99123):

i got a 60 on my test :/

OpenStudy (anonymous):

Okay so what's the answer?!